OpenStudy (anonymous):

c varies directly as d, and c = 24 when d = 6. Find d when c = 36. 1 a. 1/4 b. 4 c. 6 d. 9

OpenStudy (anonymous):

So scale factor

OpenStudy (anonymous):

\[\frac{ 24 }{ 6 } = \frac{ 36 }{ x }\]

OpenStudy (anonymous):

To solve this, you can cross multiple and set equal. 24x = 216, solve for x.

OpenStudy (anonymous):

huh

OpenStudy (anonymous):

The problem I set up is Proportions. if there is 24 of c, then there will be 6 of d.

OpenStudy (anonymous):

so what do i do?

OpenStudy (anonymous):

You would have c, which was 24, over d which was 6, but then since they say now that if c was to be 36, what would d be. so 36 over x

OpenStudy (anonymous):

Any specific points you're curious on?

OpenStudy (anonymous):

\[\frac{ 24 }{ 6 } = \frac{ 36 }{ }\]